# Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries

### v1

@@ Discussion
We combined X’s Community Notes data, the algorithmic representation learned by this service, and multi-country ideological scaling for characterizing its operation in polarized settings. We showed, in 13 countries, that notes are requested, proposed, and displayed under posts authored by accounts throughout the political spectrum. The majority of proposed Community Notes discuss politics, and predominantly rely on news outlet articles as sources, followed by X posts and Wikipedia articles.
@@
Our results highlight shortcomings of this crowd-sourced system. While notes displayed on X (i.e., having reached Helpful status) have been shown to align with professional fact-checking assessments  (?, ?), factual but polarizing content may still fail to be resolved as Helpful. Only one-tenth of proposed notes ultimately appear on X, with even lower rates for notes associated with politically prominent figures, popular accounts, and divisive issues. More generally, and by design, the most polarizing content relating to the structuring dimension of national politics typically cannot be algorithmically resolved by Community Notes. This limitation is especially relevant to events that activate such structural lines of division, such as national elections, pointing toward the potential complementarity between expert and crowd-sourced fact-checking, as originally intended by X  (?).
@@ Acknowledgments
Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries
Paul Bouchaud, Pedro Ramaciotti
[…]
@@ B.2 Posts associated with notes
The 616,943 collected posts associated with a proposed Community Notes (hereinafter refer to as “annotated posts”) were authored by 140,913 X accounts, averaging 7.8 notes per author (with median equal to 2); among accounts with at least one annotated post. Elon Musk’s account (@elonmusk) had the highest number of notes, with over nineteen thousand notes proposed by the community under more than 3,300 posts, followed by conspiracy theorist account
@BGatesIsaPyscho with over four thousand notes proposed by the community under more than 2,700 posts; we display these account profiles in Figure S8. Similarly, Elon Musk’s account (@elonmusk) had the highest number of posts (at least 3,087) with Community Notes being requested, followed by Kamala Harris’s campaign account (@KamalaHQ) with at least 1,203 posts with requests. We display in Figure S7.A the distribution per account of posts with requests for Community Notes.
(a) Profile of X account: @elonmusk.
@@ C.2 Web sources used in notes
References to x.com can be further disaggregated into user-generated content (posts, profiles, search results) and platform documentation. Specifically, 18.3% of Community Notes referencing to x.com or twitter.com refer to one of the following ressources: communitynotes.x.com,
help.x.com, blog.x.com, business.x.com, and
x.com/en/tos (combined with twitter.com variations).
@@ C.2.1 News outlets
reuters.com, apnews.com, news.yahoo.com, news.yahoo.co.jp, bbc.co.uk, nhk.or.jp, bbc.com, theguardian.com, cnn.com, nytimes.com, impress.co.jp, forbes.com, globo.com, vice.com, usatoday.com, washingtonpost.com, nbcnews.com, ft.com, dailymail.co.uk, npr.org, news.yahoo.com, timesofisrael.com,
newsweek.com, uol.com.br, cbsnews.com, afp.com, asahi.com, aljazeera.com, independent.co.uk, abcnews.go.com, infobae.com, lemonde.fr, politico.com, telegraph.co.uk, foxnews.com, wsj.com, cnbc.com, pbs.org, cbc.ca, france24.com, jpost.com, time.com, francetvinfo.fr, thehill.com, businessinsider.com, elpais.com, cnnbrasil.com.br, liberation.fr, indiatimes.com, euronews.com, elmundo.es, haaretz.com, politico.eu, estadao.com.br, leparisien.fr, bloomberg.com, lefigaro.fr, bfmtv.com, ctvnews.ca, dw.com, eldiario.es, gazetadopovo.com.br, hindustantimes.com, indiatoday.in, jiji.com, lanacion.com.ar, latimes.com, mainichi.jp, metropoles.com, ndtv.com, nikkei.com, nypost.com, spiegel.de, poder360.com.br, thehindu.com, yomiuri.co.jp, ynet.co.il, tagesschau.de, sankei.com, 20minutes.fr, 20minutos.es, aa.com.tr, abc.es, bild.de, br.de, cnn.co.jp, economist.com, elconfidencial.com, haaretz.co.il, i24news.tv, irishtimes.com, lepoint.fr, kyivindependent.com, lavanguardia.com, nationalpost.com, ndr.de, theatlantic.com, theverge.com, zeit.de, wired.com, express.co.uk, vox.com, rtve.es, rte.ie, rp.pl, publico.es,
news.sky.com, itmedia.co.jp, axios.com, aap.com.au,
veja.abril.com.br, economictimes.com, indianexpress.com, standard.co.uk, elespanol.com, clarin.com, tokyo-np.co.jp, ouest-france.fr, welt.de, voanews.com, washingtonexaminer.com, tribune.com.pk, sueddeutsche.de, faz.net, thetimes.co.uk,
dawn.com, tagesspiegel.de, news.com.au, pagina12.com.ar, elperiodico.com, israelhayom.co.il, tf1info.fr, globalnews.ca, radiofrance.fr, timesnownews.com, tvn24.pl, lasexta.com,
msnbc.com, usnews.com
@@ C.3 Topics of posts and notes
While only 11.9% of annotated political posts, having been retrieved, receive a note with Helpful Status, this rate increases to 18.5% for non-political posts; statistical significance confirmed through one-vs-rest chi-squared tests (p-value $<10^{-4}$).
@@ C.5.2 Accounts receiving most notes per country
  • United States: @elonmusk, @BGatesIsaPyscho,
    @jacksonhinklle, @iluminatibot, @EndWokeness, @KamalaHQ, @JoeBiden, @VigilantFox, @libsoftiktok, @POTUS
@@
  • United Kingdom: @Conservatives, @RishiSunak, @BladeoftheS, @PolitlcsUK, @Nigel_Farage,
    @BGatesIsaPyscho, @darrengrimes_, @Keir_Starmer, @TRobinsonNewEra, @elonmusk
  • Spain: @vitoquiles, @el_pais, @CapitanBitcoin, @eldiarioes, @elmundoes, @FonsiLoaiza,
    @sanchezcastejon, @wallstwolverine, @abc_es, @AlertaNews24
@@
  • France: @CerfiaFR, @silvano_trotta, @BFMTV,
    @JLMelenchon, @AlertesInfos, @RimaHas, @ALeaument, @f_philippot, @sandrousseau, @franceinfo
@@
  • India: @MrSinha_, @zoo_bear, @INCIndia,
    @IndianTechGuide, @dhruv_rathee, @ANI, @rishibagree, @INCKerala, @RahulGandhi, @TimesAlgebraIND
  • Argentina: @JMilei, @porqueTTarg, @therealbuni, @ElTrumpista, @laderechadiario, @infobae,
    @clarincom, @TugoNews, @GordoDan_, @TraductorTeAma
  • Israel: @YinonMagal, @netanyahu, @bezalelsm,
    @itamarbengvir, @ShaykhSulaiman, @jacksonhinklle, @Now14Israel, @TallyGotliv, @rothmar, @amit_segal,
@@ D.3 Outcomes by account
  • Over-Moderated: @BGatesIsaPyscho (33.1%),
    @iluminatibot (26.1%), @thehealthb0t (28.8%), @DailyLoud (41.4%), @Evony_TKR (78.1%), @DramaAlert (33.1%), @QTHESTORMM (52.0%), @realstewpeters (29.3%), @dom_lucre (23.0%), @tassagency_en (30.1%)
@@
  • Over-Moderated: @kirinjisinken (81.6%), @tweetsoku1 (40.0%), @geoscience16 (47.5%), @kkkfff1234k (45.1%), @hibakuyada (41.3%), @takui_aka (36.2%), @jhmdrei (30.8%), @giantearthquake (39.3%),
    @foodloss_sos (83.3%), @himuro398 (26.4%)
@@
  • Under-Moderated: @CerfiaFR (5.4%), @RimaHas (3.9%), @AlertesInfos (5.9%), @julienbahloul (0.0%), @CaronAymericoff (3.7%), @EmmanuelMacron (1.5%), @CordierAlice2 (0.0%), @Pediavenir (0.0%),
    @ArnaultRaphael (4.9%), @emma_ducros (0.0%)
@@
Across different countries, a consistent observation emerges: politically divisive accounts —whether associated with
@KamalaHarris, @tamakiyuichiro, @sanchezcastejon, @EmmanuelMacron, or @LulaOficial— display smaller fraction of post with note having reached Helpful Status compared to the general population. Conversely, entertainment accounts (@DramaAlert, @CR7Brasil) and conspiracy theory promoters (e.g., @BGatesIsaPsycho, @silvano_trotta) tend to generate high agreement and exhibit elevated moderation rates.

### v2

@@ Discussion
We combined X’s Community Notes data, the algorithmic representation learned by this service, and ideological scaling for characterizing its operation in polarized settings. We showed, in 13 countries, that notes are requested, proposed, and displayed under posts authored by accounts throughout the political spectrum. The majority of proposed Community Notes discuss politics, and predominantly rely on news outlet articles as sources, followed by X posts and Wikipedia articles.
@@
Our results highlight shortcomings of this crowd-sourced system. While notes displayed on X (i.e., having reached Helpful status) have been shown to align with professional fact-checking assessments (?, ?), factual but polarizing content may still fail to be resolved as Helpful. Only one-tenth of proposed notes ultimately appear on X, with even lower rates for notes associated with politically prominent figures, popular accounts, and divisive issues. More generally, and by design, the most polarizing content relating to the structuring dimension of national politics typically cannots be algorithmically resolved by Community Notes, regardless of their factuality. This limitation is especially relevant to events that activate such structural lines of division, such as national elections, pointing toward the potential complementarity between expert and crowd-sourced fact-checking, as originally intended by X (?).
@@
Materials and Methods
Data collection
X publicly and continuously releases all Community Notes contributions enabling third-party scrutiny. As of March 1, 2025, 1.49 million X users have submitted 4.83 million requests for Community Notes across 1.91 million posts, 1.88 millions notes were submitted and rated over 135.17 million times.
To comply with platform policies, user privacy rights, and deletion requests, X’s Community Notes dataset only discloses post IDs associated with notes, without further information on post content or author. To enable contextual analysis beyond ratings, we collected post content and author information, associated to those notes. As of March 2025, 616 104 X posts associated with 83.5% of non-deleted Community Notes were collected. The remaining posts were unavailable at the time of collection, i.e., having been removed. Additionally, out of 966k posts with at least two user requests for a Community Note, we collected 804k posts (83.2%); the remaining posts were unavailable at the time of collection in May 2025. Importantly, it should be noted that such collection process (as opposed to live stream collection or access to data reserved to the platform) introduces an inherent observational bias. Indeed, some analyses are limited to posts that remain available at collection time, which then captures the tendency of users to remove their posts after the note reached Helpful Status. This observational bias should be considered when interpreting the following results. Analyses that rely solely on the content of Community Notes, however, such as source analysis, are not marred by such limitation.
Topic Analysis
To classify community note topics, we employed a two-tiered approach. We first translated all posts to English using Facebook’s M2M100 translation model (?), ensuring consistent classification performance across original languages. For general topic classification, we used Antypas et al.’s topic taxonomy and model (?), selected for its open-weight architectures and top-tier performance. For fine-grained political classification, we employed Manifestoberta (?), a language model based on XLM-RoBERTa (?), trained on 1.6 million annotated statements from the Manifesto Corpus (?) following the Manifesto Project coding scheme (?). We report validation results for both approaches in Supplementary Information. Finally, using a curated set of keywords listed in Supplementary Information, we extracted community notes related to elections held in the United States (2024 presidential election), United Kingdom (2024 general election), France (2024 legislative election), and Germany (2025 federal elections).
Latent Ideology
We learn biases $\beta_{0},~\beta_{r},~\beta_{n}$ and latent ideologies $\theta_{r},~\theta_{n}$ for notes and raters from the publicly released community notes and ratings. We replicate X’s approach (?, ?) by implementing the matrix factorization described in Equation Community Notes undermoderate polarizing content by design creating risks in electoral processes; computational details and validation are provided in Supplementary Information.
To interpret the learned latent ideology dimension in terms of Left-Right and Anti-Elite positions, we use the ideological scaling of X users developed by Ramaciotti et al. (?). This approach extends Barberá’s (?) methodology to identify multidimensional ideological positions of users using political survey data across countries. We collected X users following Members of Parliament (MPs) across 23 countries: France, Germany, United States, United Kingdom, Spain, Brazil, Australia, Italy, Israel, Argentina, Denmark, Sweden, Canada, Poland, the Netherlands, Switzerland, South Africa, Nigeria, Finland, Mexico, New Zealand, Japan, and Belgium. We retained only users following at least three MPs and followed by at least 25 accounts, following Barberá’s protocol (?, ?, ?). We then estimate ideal positions for users and MPs using Correspondence Analysis (?) on the follower-MP adjacency matrix. Validation and interpretation are reported in Supplementary Information.
@@ Acknowledgments
The authors thank Jean-Philippe Cointet and Armin Pournaki for their comments, and Hiroki Yamashita, Jimena Royo-Letelier and Armin Pournaki for their assistant in data collection and treatment.
@@
Author contributions
P.B. and P.R. collected the data, designed the research, performed the analyses, and wrote the paper.
Competing interests
There are no competing interests to declare.
@@
Community Notes undermoderate polarizing content by design creating risks in electoral processes
Paul Bouchaud and Pedro Ramaciotti
@@ B.2 Posts associated with notes
The 616,943 collected posts associated with a proposed Community Notes (hereinafter refer to as “annotated posts”) were authored by 140,913 X accounts, averaging 7.8 notes per author (with median equal to 2); among accounts with at least one annotated post. Elon Musk’s account (@elonmusk) had the highest number of notes, with over nineteen thousand notes proposed by the community under more than 3,300 posts, followed by conspiracy theorist account @BGatesIsaPyscho with over four thousand notes proposed by the community under more than 2,700 posts; we display these account profiles in Figure S8. Similarly, Elon Musk’s account (@elonmusk) had the highest number of posts (at least 3,087) with Community Notes being requested, followed by Kamala Harris’s campaign account (@KamalaHQ) with at least 1,203 posts with requests. We display in Figure S7.A the distribution per account of posts with requests for Community Notes.
(a) Profile of X account: @elonmusk
@@ C.2 Web sources used in notes
References to x.com can be further disaggregated into user-generated content (posts, profiles, search results) and platform documentation. Specifically, 18.3% of Community Notes referencing to x.com or twitter.com refer to one of the following ressources: communitynotes.x.com, help.x.com, blog.x.com, business.x.com, and x.com/en/tos (combined with twitter.com variations).
@@ C.2.1 News outlets
reuters.com, apnews.com, news.yahoo.com, news.yahoo.co.jp, bbc.co.uk, nhk.or.jp, bbc.com, theguardian.com, cnn.com, nytimes.com, impress.co.jp, forbes.com, globo.com, vice.com, usatoday.com, washingtonpost.com, nbcnews.com, ft.com, dailymail.co.uk, npr.org, news.yahoo.com, timesofisrael.com, newsweek.com, uol.com.br, cbsnews.com, afp.com, asahi.com, aljazeera.com, independent.co.uk, abcnews.go.com, infobae.com, lemonde.fr, politico.com, telegraph.co.uk, foxnews.com, wsj.com, cnbc.com, pbs.org, cbc.ca, france24.com, jpost.com, time.com, francetvinfo.fr, thehill.com, businessinsider.com, elpais.com, cnnbrasil.com.br, liberation.fr, indiatimes.com, euronews.com, elmundo.es, haaretz.com, politico.eu, estadao.com.br, leparisien.fr, bloomberg.com, lefigaro.fr, bfmtv.com, ctvnews.ca, dw.com, eldiario.es, gazetadopovo.com.br, hindustantimes.com, indiatoday.in, jiji.com, lanacion.com.ar, latimes.com, mainichi.jp, metropoles.com, ndtv.com, nikkei.com, nypost.com, spiegel.de, poder360.com.br, thehindu.com, yomiuri.co.jp, ynet.co.il, tagesschau.de, sankei.com, 20minutes.fr, 20minutos.es, aa.com.tr, abc.es, bild.de, br.de, cnn.co.jp, economist.com, elconfidencial.com, haaretz.co.il, i24news.tv, irishtimes.com, lepoint.fr, kyivindependent.com, lavanguardia.com, nationalpost.com, ndr.de, theatlantic.com, theverge.com, zeit.de, wired.com, express.co.uk, vox.com, rtve.es, rte.ie, rp.pl, publico.es, news.sky.com, itmedia.co.jp, axios.com, aap.com.au, veja.abril.com.br, economictimes.com, indianexpress.com, standard.co.uk, elespanol.com, clarin.com, tokyo-np.co.jp, ouest-france.fr, welt.de, voanews.com, washingtonexaminer.com, tribune.com.pk, sueddeutsche.de, faz.net, thetimes.co.uk, dawn.com, tagesspiegel.de, news.com.au, pagina12.com.ar, elperiodico.com, israelhayom.co.il, tf1info.fr, globalnews.ca, radiofrance.fr, timesnownews.com, tvn24.pl, lasexta.com, msnbc.com, usnews.com
@@ C.3 Topics of posts and notes
While only 11.9% of annotated political posts, having been retrieved, receive a note with Helpful Status, this rate increases to 18.5% for non-political posts; statistical significance confirmed through one-vs-rest chi-squared tests (p-value $10^{-4}$).
@@ C.5.2 Accounts receiving most notes per country
  • United States: @elonmusk, @BGatesIsaPyscho, @jacksonhinklle, @iluminatibot, @EndWokeness, @KamalaHQ, @JoeBiden, @VigilantFox, @libsoftiktok, @POTUS
@@
  • United Kingdom: @Conservatives, @RishiSunak, @BladeoftheS, @PolitlcsUK, @Nigel_Farage, @BGatesIsaPyscho, @darrengrimes_, @Keir_Starmer, @TRobinsonNewEra, @elonmusk
  • Spain: @vitoquiles, @el_pais, @CapitanBitcoin, @eldiarioes, @elmundoes, @FonsiLoaiza, @sanchezcastejon, @wallstwolverine, @abc_es, @AlertaNews24
@@
  • France: @CerfiaFR, @silvano_trotta, @BFMTV, @JLMelenchon, @AlertesInfos, @RimaHas, @ALeaument, @f_philippot, @sandrousseau, @franceinfo
@@
  • India: @MrSinha_, @zoo_bear, @INCIndia, @IndianTechGuide, @dhruv_rathee, @ANI, @rishibagree, @INCKerala, @RahulGandhi, @TimesAlgebraIND
  • Argentina: @JMilei, @porqueTTarg, @therealbuni, @ElTrumpista, @laderechadiario, @infobae, @clarincom, @TugoNews, @GordoDan_, @TraductorTeAma
  • Israel: @YinonMagal, @netanyahu, @bezalelsm, @itamarbengvir, @ShaykhSulaiman, @jacksonhinklle, @Now14Israel, @TallyGotliv, @rothmar, @amit_segal,
@@ D.2 Outcomes by topic
Additionally, we consider the subset of community notes related to elections held in the United States (2024 presidential election), United Kingdom (2024 general election), France (2024 legislative election), and Germany (2025 federal elections). For each election, we curated the following lists of keywords (translated into English) related to the elections and political parties:
  • United States election, vote, voting, voter, ballot, campaign, Trump, Harris, Kamala, Donald Trump, Kamala Harris, presidential, president, POTUS, swing state, battleground, electoral, electoral college, Republican, Democrat, Democratic, GOP, midterm, primary, caucus, poll, polling, Biden, Joe Biden, debate, rally, MAGA, election day, early voting, mail-in ballot, absentee ballot, Pennsylvania, Michigan, Wisconsin, Arizona, Georgia, Nevada, North Carolina, electoral votes, red wave, blue wave, turnout, registered voter, Vance, JD Vance, Walz, Tim Walz, running mate, VP, vice president, inauguration, transition, concession, recount, projection, race called, swing district, Senate race, House race, congressional, gubernatorial, ballot measure, referendum, voter suppression, voter fraud, election integrity, certification, electoral count, November 5
  • United Kingdom general election, UK election, Britain election, British election, Parliament, parliamentary, House of Commons, Westminster, MP, member of parliament, constituency, seat, Labour, Labour Party, Conservative, Tory, Lib Dem, Liberal Democrats, Reform UK, Reform, Green Party, Greens, SNP, Scottish National Party, Plaid Cymru, DUP, Democratic Unionist Party, Sinn Féin, Alliance Party, UUP, Ulster Unionist Party, Keir Starmer, Rishi Sunak, Ed Davey, Nigel Farage, John Swinney, Rhun ap Iorwerth, Carla Denyer, Adrian Ramsay, Kemi Badenoch, Liz Truss, Boris Johnson, Jeremy Corbyn, Theresa May, Rachel Reeves, Angela Rayner, Prime Minister, PM, first-past-the-post, FPTP, proportional representation, PR, vote, voting, voter, ballot, polling, poll, turnout, majority, landslide, hung parliament, coalition, minority government, government formation, opposition, manifesto, campaign, election campaign, snap election, dissolution, prorogation, July 4, May 22, King Charles, monarch, 10 Downing Street, Downing Street, Number 10, swing state, swing seat, marginal seat, safe seat, red wall, blue wall, tactical voting, exit poll, by-election, Scotland, Wales, England, Northern Ireland, Scottish, Welsh, NHS, National Health Service, immigration, Brexit, cost of living, economy, tax, taxation, Rwanda, Rwanda scheme, D-Day, debate, televised debate, voter ID, photo ID, postal vote, early voting, constituency boundary, boundary review, redistricting, declaration, count, recount, incumbent, challenger, cabinet, shadow cabinet, minister, secretary of state, backbencher, frontbencher, whip, Speaker, Chancellor, Home Secretary, Foreign Secretary, election 2024
  • France: legislative election, French election, election, National Assembly, deputies, RN, National Rally, Marine Le Pen, Jordan Bardella, Emmanuel Macron, NFP, New Popular Front, Popular Front, LFI, France Insoumise, Jean-Luc Mélenchon, PS, Socialist Party, Olivier Faure, Ensemble, Renaissance, Gabriel Attal, Michel Barnier, François Bayrou, Élisabeth Borne, LR, Republicans, PCF, Communist Party, EELV, Ecologists, Greens, Marine Tondelier, Horizons, MoDem, snap election, dissolution, cohabitation, hung parliament, first round, second round, runoff, three-way race, four-way race, June 30, July 7, turnout, participation, constituency, vote, voting, far-right, radical right, left-wing, coalition, alliance, government formation, Prime Minister, President, parliament, majority, absolute majority, minority government, opposition, projections, polls, 2024 election
  • Germany: German federal election, Bundestag, German election, CDU, CSU, CDU/CSU, SPD, AfD, Greens, FDP, The Left, BSW, Sahra Wagenknecht, Friedrich Merz, Olaf Scholz, Alice Weidel, Christian Lindner, traffic light coalition, grand coalition, coalition, chancellor, vote, voting, first vote, second vote, constituency, Federal Returning Officer, confidence vote, snap election, early election, February 23, voter turnout, government, opposition, Federal President, Frank-Walter Steinmeier, Bavaria, Baden-Württemberg, North Rhine-Westphalia, Saxony, Thuringia, Brandenburg, Saxony-Anhalt, Mecklenburg-Western Pomerania, Berlin, Hamburg, Hesse, Lower Saxony, Rhineland-Palatinate, Saarland, Schleswig-Holstein, Bremen, East Germany, West Germany, right-wing populism, right-wing, far-right, firewall, proportional representation, electoral threshold, five percent threshold, 5% threshold, overhang mandate, balance seat, direct mandate, state list, party list, manifesto, electoral program, coalition negotiations, coalition talks, government formation, minority government, election 2025
For robustness, we bootstrap over keywords by selecting community notes containing a random sample of 50% of keywords, measuring the fraction reaching helpful status, and reporting the average fraction and 95% confidence interval over 250 repetitions.
@@ D.3 Outcomes by account
  • Over-Moderated: @BGatesIsaPyscho (33.1%), @iluminatibot (26.1%), @thehealthb0t (28.8%), @DailyLoud (41.4%), @Evony_TKR (78.1%), @DramaAlert (33.1%), @QTHESTORMM (52.0%), @realstewpeters (29.3%), @dom_lucre (23.0%), @tassagency_en (30.1%)
@@
  • Over-Moderated: @kirinjisinken (81.6%), @tweetsoku1 (40.0%), @geoscience16 (47.5%), @kkkfff1234k (45.1%), @hibakuyada (41.3%), @takui_aka (36.2%), @jhmdrei (30.8%), @giantearthquake (39.3%), @foodloss_sos (83.3%), @himuro398 (26.4%)
@@
  • Under-Moderated: @CerfiaFR (5.4%), @RimaHas (3.9%), @AlertesInfos (5.9%), @julienbahloul (0.0%), @CaronAymericoff (3.7%), @EmmanuelMacron (1.5%), @CordierAlice2 (0.0%), @Pediavenir (0.0%), @ArnaultRaphael (4.9%), @emma_ducros (0.0%)
@@
Across different countries, a consistent observation emerges: politically divisive accounts —whether associated with @KamalaHarris, @tamakiyuichiro, @sanchezcastejon, @EmmanuelMacron, or @LulaOficial— display smaller fraction of post with note having reached Helpful Status compared to the general population. Conversely, entertainment accounts (@DramaAlert, @CR7Brasil) and conspiracy theory promoters (e.g., @BGatesIsaPsycho, @silvano_trotta) tend to generate high agreement and exhibit elevated moderation rates.